# Nemotron-4 15B Technical Report

### 26 Feb 2024 v1

@@ 1   Introduction
As demonstrated in Figure 1, Nemotron-4 15B exhibits high downstream accuracies across a wide range of English, code, and multilingual evaluation areas. In comparison to leading similarly-sized, open models we show that Nemotron-4 15B is significantly better than LLaMA-2 34B (Touvron et al., 2023b), which has over twice the number of parameters, and is better than Mistral 7B (Jiang et al., 2023) on all English evaluation areas. Additionally, Nemotron-4 15B achieves competitive accuracies to QWEN 14B (Bai et al., 2023) and Gemma 7B (Gemma Team, 2024). In a comparison across a wide range of programming languages, we find that Nemotron-4 15B achieves better accuracy overall, and in particular on low-resource programming languages, than Starcoder (Li et al., 2023), a code-specific model, and Mistral 7B. As Nemotron-4 15B was trained on significant amount of multilingual data, it is currently the state-of-the-art general purpose model in its size class on all multilingual benchmarks. We find that Nemotron-4 is better than PALM 62B-Cont (Slav Petrov and et al., 2023), and also outperforms multilingual-specific models such as XGLM (Lin et al., 2022) and mGPT (Shliazhko et al., 2022).
@@ 2   Architecture Details
Nemotron-4 uses a standard decoder-only Transformer architecture (Vaswani et al., 2017), with causal attention masks. Exact hyper-parameters affecting size are shown in Table 1. Nemotron-4 has 3.2 billion embedding parameters and 12.5 billion non-embedding parameters. We use Rotary Position Embeddings (RoPE) (Su et al., 2021), SentencePiece tokenizer (Kudo and Richardson, 2018), squared ReLU activations in the MLP layers, no bias terms, dropout of 0, and untied input-output embeddings. We use Grouped Query Attention (GQA) (Ainslie et al., 2023) for faster inference and lower memory footprint.
@@
The English data corpus consists of curated documents from a variety of sources and domains including web documents, news articles, scientific papers, books, etc and the distribution used in our pre-training set is highlighted in Figure 2. The code and multilingual data consists of a diverse set of natural and programming languages. We find that appropriately sampling tokens from these languages is key to strong accuracies in these domains. We share the distributions used for both code and multilingual tokens in our pre-training dataset in Figure 3 and Figure 4 respectively.
Figure 3: Data distribution of coding languages used for pre-training. The number within each bar indicates the percent of the overall code distribution that an individual language comprises.
@@
Figure 4: Data distribution of multilingual languages used for pre-training. The number within each bar indicates the percent of the overall multilingual distribution that an individual language comprises.
@@ 3   Results
Table 3: Result on standard reasoning benchmarks in the zero-shot setting. We report the average across all tasks where possible for a fair comparison. The values marked with *** are read from  Gemma Team (2024)
@@ 4   Conclusion
We present Nemotron-4 15B, a decoder-only transformer-based large language model. It is trained on 8 trillion tokens spanning English, 53 additional natural languages as well as 43 programming languages. Nemotron-4 15B exhibits the strongest multilingual performance of any general purpose language model at its scale – even outperforming models specialized for the multilingual domain. Nemotron-4 demonstrates that large language models can be pre-trained on more tokens than previously estimated with exceptional results.

### 27 Feb 2024 v2

@@ 1   Introduction
As demonstrated in Figure 1, Nemotron-4 15B exhibits high downstream accuracies across a wide range of English, code, and multilingual evaluation areas. In comparison to leading similarly-sized, open models we show that Nemotron-4 15B is significantly better than LLaMA-2 34B (Touvron et al., 2023b), which has over twice the number of parameters, and is better than Mistral 7B (Jiang et al., 2023) on all English evaluation areas. Additionally, Nemotron-4 15B achieves competitive accuracies to QWEN 14B (Bai et al., 2023) and Gemma 7B (Gemma Team, 2024). In a comparison across a wide range of programming languages, we find that Nemotron-4 15B achieves better average accuracy, and in particular on low-resource programming languages, than Starcoder (Li et al., 2023), a code-specific model, and Mistral 7B. As Nemotron-4 15B was trained on significant amount of multilingual data, it is currently the state-of-the-art general purpose model in its size class on all multilingual benchmarks. We find that Nemotron-4 is better than PALM 62B-Cont (Slav Petrov and et al., 2023), and also outperforms multilingual-specific models such as XGLM (Lin et al., 2022) and mGPT (Shliazhko et al., 2022).
@@ 2   Architecture Details
Nemotron-4 uses a standard decoder-only Transformer architecture (Vaswani et al., 2017), with causal attention masks. Exact hyper-parameters affecting size are shown in Table 1. Nemotron-4 has 3.2 billion embedding parameters and 12.5 billion non-embedding parameters. We use Rotary Position Embeddings (RoPE) (Su et al., 2021), SentencePiece tokenizer (Kudo and Richardson, 2018), squared ReLU activations in the MLP layers, no bias terms, dropout rate of zero, and untied input-output embeddings. We use grouped query attention (GQA) (Ainslie et al., 2023) for faster inference and lower memory footprint.
@@
The English corpus consists of curated documents from a variety of sources and domains including web documents, news articles, scientific papers, books, etc and the distribution used in our pre-training set is highlighted in Figure 2. The code and multilingual data consists of a diverse set of natural and programming languages. We find that appropriately sampling tokens from these languages is key to strong accuracies in these domains. We share the distributions used for both code and multilingual tokens in our pre-training dataset in Figure 3 and Figure 4 respectively.
Figure 3: Data distribution of the 43 programming languages used for pre-training. The number within each bar indicates the percent of the overall code distribution that an individual language comprises.
@@
Figure 4: Data distribution of the 53 natural languages, aside from English,we used for pre-training. The number within each bar indicates the percent of the overall multilingual distribution that an individual language comprises.
@@ 3.1 Commonsense Reasoning
Table 3: Results on standard reasoning benchmarks in the zero-shot setting. We report the average across all tasks where possible for a fair comparison. The values marked with *** are read from  Gemma Team (2024)
@@ 4   Conclusion
We present Nemotron-4 15B, a decoder-only transformer-based large language model. It is trained on 8 trillion tokens spanning English, 53 additional natural languages as well as 43 programming languages. Nemotron-4 15B exhibits the strongest multilingual performance of any general purpose language model at its scale – even outperforming models specialized for the multilingual domain. Nemotron-4 demonstrates that pre-training sets for large language models can continue to be scaled up even further in order to improve the abilities of models.